ABCD is a quadrilateral with coordinates A (-5,2), B (6,2), C (1,-5), and D (-2,-3). Find the verticals of A'B'C'D under the translation (X+3,y - 6).


1
Expert's answer
2021-12-13T07:20:09-0500

translation (X+3,y - 6) has translation vector "\vec{a} (3, -6)" , so each point shifts 3 points to the right and 6 points downward, which means

A' = (-5 + 3, 2 - 6) = (-2, -4)

B' = (6 + 3, 2 - 6) = (9, -4)

C' = (1 + 3, -5 - 6) = (4, -11)

D' = (-2 + 3, -3 - 6) = (1, -9)